8x+5y=-37
6x-5y=-89

Answer by alicealc(293)

use the elimination procedure:
8x + 5y = -37
6x - 5y = -89
--------------- (+)
14x = -126
x = -126/14 = -9

8x + 5y = -37
8*(-9) + 5y = -37
-72 + 5y = -37
5y = -37 + 72
5y = 35
y = 35/5 = 7

so, the solution is x = -9, y = 7